2m + n = 0
m + 2n = 3

Answer by Alan3354(69443)

2m + n = 0
m + 2n = 3
--------------
If you want to solve for m & n, you should say that.
================
2m + n = 0
m + 2n = 3
------
1 method is substitution:
n = -2m
Sub for n in the 2nd eqn
m + 2*(-2m) = 3
m - 4m = 3
-3m = 3
m = -1
----
n = -2m = +2
======================
Another method is elimination:
2m + n = 0
m + 2n = 3
---
Multiply the 1st eqn by 2:
4m + 2n = 0
m + 2n = 3
--------------------- Subtract
3m = -3
m = -1
etc